Two charges are separated by a distance of 0.03 meters. Identify the magnitude of the net

force exerted by each charge on one another.


q1 = +25 × 10^−9C and q2 = −75 × 10^−9C


1
Expert's answer
2021-01-23T10:18:40-0500

The Coulomb's law says

"F=k\frac{|q_1||q_2|}{r^2}""F=9\times 10^9\times\frac{25\times 10^{-9}\times 75\times 10^{-9} }{0.03^2}=0.019\:\rm N"